I'm making an ern for My Uncle's ashes. He was a Forester. I need to know how big is a sphere 200 CUBIC INCHES in a stump that is 8w.8d.6h


Answer by checkley77(12844)

V=4/3PIR^3
300=1.333*3.14*R^3
300=4.1856*R^3
R^3=300/4.1856
R^3=71.67
R=CUBERT71.67
R=4.154 INCHES FOR THE 300 CUBIC INCH SPHERE.